OpenStudy (anonymous):

For what values of n is y=x^n a solution of the differential equation \[x^2y\prime\prime+xy\prime-y=0\]

OpenStudy (anonymous):

\[ x^2 y''(x)+x y'(x)-y(x)=0\\ (n-1) n x^n+n x^n-x^n=0\\ (n-1) (n+1) x^n=0\\ n=1\\ n=-1 \]